The triangle has height (x+1)cm and base (2x-3) cm. It's area is 9 cm squared.
how do you show that 2x squared - x -21 = 0

Answer by mananth(16946)


Area of triangle = 1/2 * base * height
Area = 1/2 *(x+1)(2x-3)=9
(x+1)(2x-3)=18
2x^2-x-3=18
2x^2-x-21=0
...
m.ananth@hotmail.ca